OpenStudy (anonymous):

Solve for x in the proportion:

OpenStudy (anonymous):

\[\frac{ 3x }{ 3x - 36 }\] = \[\frac{ -1 }{ x }\]

OpenStudy (mertsj):

Did you cross multiply?

OpenStudy (mertsj):

Also it would make your life easier if you would divide numerator and denominator of the first fraction by 3 before you cross multiply.

OpenStudy (anonymous):

\[\frac{ x }{ x - 12 }\] ? @Mertsj

OpenStudy (mertsj):

Yes. So you you have:

OpenStudy (mertsj):

\[\frac{x}{x-12}=\frac{-1}{x}\]

OpenStudy (mertsj):

Cross multiply and solve.

OpenStudy (anonymous):

Alrighty, one second.

OpenStudy (anonymous):

x^2 = -x + 12 ?

OpenStudy (mertsj):

Now add x to both sides and subtract 12 from both sides so that the right side will be 0

OpenStudy (anonymous):

12 = x^2 + x

OpenStudy (anonymous):

0 = x^2 + x - 12 how do I solve?

OpenStudy (mertsj):

factor x^2+x-12

OpenStudy (anonymous):

Oh wow, thanks (:

OpenStudy (mertsj):

What did you get for your answers?

OpenStudy (anonymous):

x = -3 and x = 4

OpenStudy (mertsj):

I was afraid of that.

OpenStudy (mertsj):